(5.2.19) Assume that random guesses are made for 8 ​multiple-choice questions on a test with 5 choices for each​ question, so that there are n=8 ​trials, each with probability of success​ (correct) given by p=20. Find the probability of no correct answers.
The probability of no correct answers is?
1
Expert's answer
2020-11-30T10:41:27-0500

Solution:

The probability of success: "p=0.2"

"P(X=0)=_0C_8 \times p^0 \times q^8=1\times1\times0.8^8=0.1678"


Answer: Probability of no correct answers is 0.1678